OpenStudy (anonymous):

Solve for the unknown measure of A in 31^2 = 27^2 + 22^2 – 2(27)(22) cos A.

OpenStudy (science0229):

First, isolate cos A by itself. Subtracted both sides by (27^2 + 22^2) and divide by 2(27)(22). Use calculator to get cos A=7/33

OpenStudy (science0229):

Correction: dvide both sides by -2(27)(22)

OpenStudy (science0229):

Now, I'm thinking that this is from a triangle using Law of Cosine. Then, angle A has to be between 0 and pi/2.

OpenStudy (science0229):

Now, solve cos A=7/33 over 0 and pi/2 to get A=arccos(7/33) (Unit is radians) Decimal form is 1.3577051269... radians. In degree, A is about 77.75331027 degrees.

OpenStudy (loser66):

To me, the number of the left hand side forms the form of (27-22)^2 = 5^2 =25 therefore, cos A = 31^2/25 = 38.44 --> no solution because cos of angle cannot be > 1

OpenStudy (loser66):

oh sorry, right hand side, not left hand side @ybarrap what do you think?

OpenStudy (science0229):

Now, 27^2 + 22^2 – 2(27)(22) cos A is not (27-22)^2

OpenStudy (science0229):

because there is cos A multiplied to -2(27)(22) If it were 27^2+22^2-2(27)(22), then you would be right.

OpenStudy (loser66):

oh yea, you are right, I am sorry. misread the problem.

OpenStudy (loser66):

My bad. :) I am sorry

OpenStudy (science0229):

That's alright. Sometimes, I make stupid mistakes, too.

OpenStudy (ybarrap):

Here's just another way to look at this problem... $$ a\bullet b=|a||b|\cos A=31^2 = 27^2 + 22^2 – 2(27)(22) \cos A\\ \implies \\ a=27\\ b=22\\ \implies A=\cos^{-1}\cfrac{31^2}{27\times22} $$

OpenStudy (science0229):

That's another way to look at it.

OpenStudy (loser66):

@ybarrap I don't get!!

OpenStudy (science0229):

He used the "vector version" of the Law of Cosine

OpenStudy (ybarrap):



OpenStudy (ybarrap):

http://en.wikipedia.org/wiki/Law_of_cosines

OpenStudy (loser66):

got it, thanks.

OpenStudy (loser66):

Master is master, junior is junior. hehehe. wait for me , Master

OpenStudy (ybarrap):

I just happened to see the form, not always obvious

OpenStudy (ybarrap):

I learn so much from you @Loser66 and everyone else here, even these problems give me practice and it reinforces my knowledge